The economy of Tuland produces only two products

books books

and

dvds dvds.

The following information is available for production and prices of​ Tuland's products for the years 2009 and 2010.

2009

2010

Quantity of books

50

55

Quantity of dvds

2

4

Price of books

​$1.00

​$1.05

Price of dvds

​$50.00

​$30.00

Using above​ information, calculate the following values.

Real GDP for 2009 using 2009 as base year equals

​$.

​(Enter your response rounded to two decimal​ places.)

Real GDP for 2009 using 2010 as base year equals

​$.

​(Enter your response rounded to two decimal​ places.)

Real GDP for 2010 using 2009 as base year equals

​$.

​(Enter your response rounded to two decimal​ places.)

Real GDP for 2010 using 2010 as base year equals

​$.

​(Enter your response rounded to two decimal​ places.)

GDP growth rate using 2009 as base year equals

​%.

​(Enter your response rounded to two decimal​ places.)

GDP growth rate using 2010 as base year equals

​%.

​(Enter your response rounded to two decimal​ places.)

The arithmetic average of the two growth rates equals

​%.

​(Enter your response rounded to two decimal​ places.)
Answer #1

Solution:-

(A). GDP for 2009 in 2009's Prices = (50 * 1) + (2 * 50)

                                                         = 50 + 100

                                                         = $150

(B). GDP for 2009 in 2010's Prices = (50 * 1.05) + (2 * 30)

                                                         = 52.5 + 60

                                                         = $112.5

(C). GDP for 2010 in 2009's Prices = (55 * 1) + (4 * 50)

                                                         = 55 + 200

                                                         = $255

(D). GDP for 2010 in 2010's Prices = (55 * 1.05) + (4 * 30)

                                                         = 57.75 + 120

                                                         = $177.75

GDP growth rate using 2009 as base year = ($255 -150) / 150*100

                                                                    = 0.7 * 100

                                                                    = 70%

GDP growth rate using 2010 as base year = (177.75 – 112.5) / 112.5 * 100

                                                                    = 0.58 * 100

                                                                    = 58%

The arithmetic average is (58 + 70) / 2 = 64%